BLUE-DETUNED PHOTOASSOCIATION SPECTRUM IN Rb_{2}

M. A. BELLOS, D. RAHMLOW, \underline{R. CAROLLO}, J. BANERJEE, E. E. EYLER, P. L. GOULD, and W. C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06269.

We report on the observation of blue-detuned photoassociation as proposed in [1] and references therein. “Blue-detuned” refers to the location of vibrational levels — energetically above the corresponding atomic asymptote. Ultracold ^{85}Rb atoms in a MOT were photoassociated to levels of the 1^{3}Π_{g} state a few hundred wavenumbers above the 5S +5P_{3/2} limit. These transitions were found to be strong even though they occur at short internuclear separations (R_{e}=10 a_{0}). Levels of the 1^{3}Π_{g} state spontaneously decay to the a^{3}Σ_{u}^{+} state, where they are detected by resonantly enhanced multiphoton ionization with time-of-flight spectroscopy. We have observed most vibrational levels of the 1^{3}Π_{g} state belonging to all of its spin-orbit components (0_{g}^{+}, 0_{g}^{−}, 1_{g}, 2_{g}). Recent unpublished ab-initio calculations [2] of these potentials show good agreement with the observed vibrational and rotational constants. This work is supported by the NSF, AFOSR, and UConn Research Foundation.

[1] M.-L. Almazor et. al., Eur. Phys. J. D 15 355 (2001)

[2] O. Dulieu, private communication